OpenStudy (anonymous):

4(x – 3) = 7x + 24 A. x = –36 B. x = 12 C. x = 4 D. x = –12

OpenStudy (anonymous):

4x-12=7x+24 4x-7x=24+12 -3x=36 -x=12 x=-12

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yup np, but do you understand it?

OpenStudy (anonymous):

Yeah (:

OpenStudy (anonymous):

what grade r u in?

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

Can you help with this one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

do you know much about exponents?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

ok hold on one sec

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

it seems that a = 3

OpenStudy (anonymous):

3^-3=1/27

OpenStudy (anonymous):

So, what's the answer?

OpenStudy (anonymous):

haha yah

OpenStudy (anonymous):

a=3

OpenStudy (anonymous):

3 is the answer?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thank you so so much!

OpenStudy (anonymous):

anymore?

OpenStudy (anonymous):

alright frosted good luck with school this year!!!!!

OpenStudy (anonymous):

Thanks! If I have anymore, I'll tell you!

OpenStudy (anonymous):

okidoki

OpenStudy (anonymous):

What is the solution to the inequality? 4x + 7 > –13 A. x < –5 B. x > –5 C. x < 5 D. x > 5 @CALCULASMASTER_1,2,3

OpenStudy (anonymous):

Let, 4x+7>-13..... 4x>-13-7 4x>20 x>5 Ans. x>5!

OpenStudy (anonymous):

Thanks so much oh my goodness!

OpenStudy (anonymous):

just let me know when you need help, Im here anytime!.....

OpenStudy (anonymous):

Which property is shown? –7 + 24 = 24 + (–7) A. opposite of a sum property B. commutative property C. distributive property D. associative property @CALCULASMASTER_1,2,3

OpenStudy (anonymous):

Please help.

OpenStudy (anonymous):

yes I am, just one sec, let me review on properties, they were such a long time ago, one sec

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

its communicative

OpenStudy (anonymous):

typo lol

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

What is the value of the expression x + (–y) when x = –3.8 and y = 1.7? A. –5.5 B. –2.1 C. 2.1 D. 5.5 This one? @CALCULASMASTER_1,2,3

OpenStudy (anonymous):

x+(-y)=x-y=(x)-(y)=(-3.8)-(1.7)=(-3.8-1.7)=-5.5.

OpenStudy (anonymous):

Any more questions?

OpenStudy (anonymous):

Help please?

OpenStudy (anonymous):

Would you like to understand how I get the answer?

OpenStudy (anonymous):

No, I understand it.

OpenStudy (anonymous):

ok, hold on

OpenStudy (anonymous):

Do you have the answer yet?

OpenStudy (anonymous):

I'm timed.

OpenStudy (anonymous):

ypu need to find 2/3 of 3/4

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

2/3times2/4=1/2

OpenStudy (anonymous):

So, 1/2 is the answer?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Woah, I am stuck on one! Help please?

OpenStudy (anonymous):

@CALCULASMASTER_1,2,3

OpenStudy (anonymous):

30

OpenStudy (anonymous):

@CALCULASMASTER_1,2,3 What is the value of the expression 15 + 3 • 4 – 8 ÷ 2? A. 17 B. 23 C. 32 D. 68 The answer is 23, correct?

OpenStudy (anonymous):

Am I right @CALCULASMASTER_1,2,3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Last one.

OpenStudy (anonymous):

Help, I'm running out of time.

OpenStudy (anonymous):

I have legit 5 minutes.

OpenStudy (anonymous):

31/15

OpenStudy (anonymous):

2 and 1/15

OpenStudy (anonymous):

Is that it???